# Solving Modular Linear Systems with a Constraint by parallel decomposition of the Smith form and extended Euclidean division modulo powers of primes divisors

### v3

@@ 3 Solution modulo $q^{r}$ where $q=p^{d}$
    From the definition of $w_{0}$, $w_{1}$ in (7), if $<w_{0},x_{0}>\mbox{ \rm{mod} }p=0$, $w_{1}\mbox{ \rm{mod} }p\neq 0$.
@@
    To construct a solution $x_{1}$, the component defining the solutions is chosen as $x_{1}=0$ when $<w_{0},x_{0}>\mbox{ \rm{mod} }p\neq 0$ and $<w_{1},x_{1}>\mbox{ \rm{mod} }\neq 0$ when $<w_{0},x_{0}>\mbox{ \rm{mod} }p=0$.
Proof follows from the steps in Lemma (2) and Theorem (2) and from the fact that $x_{1}$ exists when $w_{1}\mbox{ \rm{mod} }p\neq 0$ such that $<w_{1},x_{1}>\mbox{ \rm{mod} }p\neq 0$. Since there is only one prime divisor $p$ in this special case of $n=q^{r}$ and $q=p^{d}$ the construction of $x_{1}$ using CRT is not necessary as in the general case.
@@ 3.1 Decomposition modulo $p^{r}$ and solution for general modulus $n$
The Problem 1 has a solution iff for each prime divisor $p$ of $n$ the modular system of Problem 2 has a solution $x(p)$ for each $p$ and $q=p$, $r=r(p)$ such that $x(p)$ satisfies $<w,x(p)>\mbox{ \rm{mod} }p\neq 0$. If $x(p)$ is a solution of the Problem 2 for each of the prime divisors $p$ of $n$ satisfying the constraint $<w,x(p)>\mbox{ \rm{mod} }p\neq 0$ and
@@
[…]
This theorem shows the decomposition of the problem from modulus $n$ to the smaller moduli $p^{r(p)}$ for prime divisors $p$ of $n$. Such a decomposition is computationally advantageous as we show next because the Smith form computation of $[A,-p^{r(p)}I]$ does not involve computation of GCDs of elements of $A$ along with $n$ but only involves division of $p^{r}$ to find the largest divisor. Hence solving the decomposed linear systems modulo $p^{r(p)}$ creates a notion of relative invariant factors which are invariant factors of $[A,-p^{r(p)}I]$. These can be computed much efficiently because the prime $p$ is known (hence also all the divisors of $p^{r}$) and hence analogous computation of the GCD of rows and columns of $A$ required in Smith form computation of $A$ are not required.
3.2 Smith form reduction of $[A,-p^{r}I]$
The above theorem gives solution of the special case of the modular system with constraint defined in Problem 2 for $q=p$. Hence it is desirable to simplify the algorithm for computation of the invariant factors of $[A,-p^{r}I]$ since all of these are powers of the prime $p$.
@@ 3.2.1 Solution of the Bezout identity with a prime power dividend
The Smith form reduction of a matrix such as $[A,-p^{r}I]$ with $p$ a prime by unimodular transformations, is based on computation of integral solutions $x_{i},y$ to the Bezout identities such as
@@
$d=r-g$, where we now consider $a_{i}$, $x_{i}$, $y$ to be in the modular ring $\mathbb{Z}/(p^{d})$. Consider a solution in the simplest case
@@
are the $p$-adic representations of $a$, $x$ and $y$ in $\mathbb{Z}/(p^{d})$ with $a_{i},x_{i},y_{i}$ in $[0,p-1]$. For any number $z$ in $\mathbb{Z}/(p^{d})$ denote $(z)=z\mbox{ \rm{mod} }p$ and $[z]=z\mbox{ \rm{div} }p$. when $(a,p^{d})$ are coprime. When $(a,p^{d})$ are coprime $a$ has no divisor $p^{k}$ for $0\leq k\leq d$.
@@
then $xb$, $yb$ give solutions of the Bezout identity for right hand side $b$. The unique solutions $x_{i}$, $y_{i}$ of co-efficients of solutions $x,y$ of (10) modulo $p^{d}$ are given by the following iterations
$\begin{array}[]{lclcl}t_{0}&:=&(a_{0}x_{0})\\ t_{0}&=&1\mbox{ \rm{mod} }p\\ \mbox{for $k=1,2,\dots,(d-1)$}&&\\ t_{k}&:=&[t_{(k-1)}]+(a_{0}x_{k}+a_{1}x_{(k-1)}+\ldots+a_{k}x_{0})\\ t_{k}&=&0\mbox{ \rm{mod} }p\\ \mbox{for $j=0,1,\ldots,(d-1)$}&&\\ t_{(d+j)}&:=&[t_{(d-1+j)}]+a_{1}x_{(d-1+j)}+a_{2}x_{(d-2+j)}+\ldots+a_{(d-1)}x_{(1+j)}\\ t_{(d+j)}+y_{j}&=&0\mbox{ \rm{mod} }p\end{array}$ (12)
@@
Proof follows by substitution of $p$-adic representations of $a$, $x$, $y$ in (10) and comparing co-efficients of powers of $p$. While each co-efficient is computed modulo $p$, the carry part which is quotient by $p$ is added to the higher power of $p$ shown as $[t_{(k-1)}]$ in the next value of $t_{k}$ in the iteration. Note that each equation $t_{k}=0$ for $k>0$ has the unknown $x_{k}$, is of the form $a_{0}x_{k}=b_{k}\mbox{ \rm{mod} }p$ where $b_{k}$ depends on previously computed $x_{0},\ldots,x_{(k-1)}$ and only inversion modulo $p$ of $a_{0}$ is required for solving all $x_{k}$. Since $x,y$ were assumed with representations in $\mathbb{Z}/(p^{d})$ the solution $x,y$ of the iterations (12) satisfy
@@
for $x$ modulo $p^{d}$. Once $x$ is solved, $ax=1+py$ for some $y$ which can be obtained by dividing $ax-1$ by $p^{d}$. Since $a$ is assumed to be in $\mathbb{Z}/(p^{d})$ for any other integer $a$ the Bezout identity solutions $x,y$ can be found by solving $(a\mbox{ \rm{mod} }p^{d})x=1$ modulo $p^{d}$.
@@
$[a,p^{d}]Q=[1,0]$
Also $[ab,p^{d}b]Q=[b,0]$. We can thus formally state
@@
If $\mbox{GCD}(a,p^{r})=b$ then there exists a unimodular matrix $Q$ such that $[a,p^{r}]Q=b$.
@@
Note that the unimodular matrix $Q$ is obtained using $x,y$ the solutions of the Bezout identity (10) for the pair $(ap^{-g},p^{d})$ for $d=r-g$ which is coprime. This computation is performed as shown in iterations (12) in Lemma 3 which involves division operations by $p$ to compute the co-efficients mod $p$ and the carry as quotient by $p$. The operations required to compute $x$, $y$ are 1) one mod $p$ inversion, 2) maximum of $d^{2}$ multiplications modulo $p$ and 3) $d$ integral divisions by $p$. The overall cost of computations of iterations (12) can be shown to be $O(l(p)d^{2})$. In general the cost of computation of the extended GCD of two integers $a,b$ is known to be approximately $O(l(a)l(b))$ where $l(a)$ is the bit length of $a$ [8]. If $a,b$ have prime factorizations
@@ 3.2.2 Solution of Bezout identity mod $p^{r}$ with multiple elements
Let $GCD(a_{1},a_{2},\ldots,a_{N},p^{r})=p^{g}$. Then there exists a unimodular matrix $Q$ such that $[a_{1},a_{2},\ldots,a_{n},p^{r}]Q=[p^{g},0,\ldots,0]$. The first column of $Q$ is $[x_{1},x_{2},\ldots,x_{N},y]^{T}$ where $x_{i}$ and $y$ are modulo $p^{r}$ and satsfy
@@
Let $Q_{1}$ be a $2$ unimodular matrix as computed in Lemma 3 such that $[a_{N},p^{r}]=[p^{t_{1}},0]$ where $p^{t_{1}}$ is the GCD of $(a_{N},p^{r})$. By induction step assume that the GCD of $(a_{(N-1)},\ldots,a_{1},p^{r})$ be $p^{t_{(N-1)}}$ and let $Q_{(N-1)}$ be a $n\times n$ unimodular matrix such that
$[a_{(N-1)},\ldots,a_{1},p^{r}]Q=[p^{(t_{(N-1)}},0,\ldots,0]$
The GCD of $(a_{N},p^{t_{(N-1)}})$ is then $p^{g}$. Let $Q^{\prime}$ be a unimodular matrix as derived in lemma 4 such that $[a_{N},p^{t_{(N-1)}}Q^{\prime}=[p^{g},0]$. Define
@@
The construction of unimodular matrix for solving the Bezout identity modulo $p^{r}$ for multiple elements $(a_{1},\ldots,a_{N})$ and GCD $p^{g}$ with $p^{r}$ uses the computations of Lemma 3 which are mod $p$ and quotients by $p$. Hence for a known prime $p$ these computations are much reduced compared to GCD computation of general $(N+1)$ elements and further solution of the general Bezout identity. In the next proposition we derive the unmodular matrix to transform the matrix $[A,-p^{r}I]$ to its Smith form.
@@
First bring the matrix $\hat{A}=[A,-p^{r}I]$ by column (unimodular) transformations $\hat{A}V_{1}$ to the upper triangular form $[p^{r}I,A]$. Let $t\leq r$ be the largest power such that $p^{t}$ divides all elements in the first row of $A$. Then $p^{t}$ is the GCD of elements in the first row of $\hat{A}V_{1}$. Hence there exists a unimodular matrix $V_{2}$ obtained from the extended Euclidean divison representation of the first row such that $\hat{A}_{2}=\hat{A}V_{1}V_{2}$ has $p^{t}$ as $(1,1)$ element while the entire first row is zero. Repeat the same process to make the first column of $\hat{A}_{2}$ equal to zero by constructing a unimodular matrix $P$ and computing $P\hat{A}_{2}$. Repeat both steps of operation by left and right unimodular matrices till $\hat{A}$ is brought to the form
@@
Since the original matrix $\hat{A}$ has some of its $2\times 2$ minors of the form $p^{2r}$ the GCD of elements of $A_{1}$ must have a divisor $p^{k_{2}}$ where $k_{2}\geq k_{1}$. Now repeat the same steps to make $A_{1}$ successively into diagonal form by unimodular row and column operations. Hence after repeating such steps the matrix $[A,-p^{r}]$ is transformed into a diagonal form with diagonal entries $p^{k_{i}}$. Next add all columns from 2nd to last to the first column and then reduce the matrix to $p^{r_{1}}$ as $(1,1)$ entry and and another matrix $A_{1}$ as above. Now we can see that $0\leq r_{1}$ and $p^{r_{1}}$ divides all other powers $p^{k_{i}}$ for $i\geq 2$. Repeating the same process now on the resultant matrix gives the Smith form of $[A,-p^{r}I]$ as claimed. ∎
@@
The reduction to Smith form of $[A,-p^{r}I]$ shown in the above Proposition uses division by $p$ in all the steps to determine largest powers of $p$ as the gcd of entries of the matrices at each step. If any element in the matrix is not divisible by $p$ then the corresponding invariant factor is $1$. Using the well known result on high probability of coprimeness of random integers it follows that the invariant factors for lower indices are highly likely to be $1$.
@@
Computing the GCD and solving the Bezout identity of a set of $N$ integers and $p^{r}$ modulo $p^{r}$ has complexity equivalent to the $O(Nd^{2}(\log p)^{2})$ where $d=r-t$ and $p^{t}$ is the GCD of $(a_{1},\ldots,a_{N},p^{r})$. For a given prime $p$ the unimodular matrix $Q$ which can transform $[a_{1},a_{2},\ldots,a_{N},p^{r}]$ as
@@ 3.3 High level algorithm to solve the modular system
4:     Form the augmented matrix $[A,-p^{r}I]$.
5:     Find unimodular integral matrices $P$, $Q$ from the procedure described in Lemma 2, to determine the Smith form $[S_{p},0]$ of the augmented matrix.

### v4

@@ 3 Solution modulo $q^{r}$ where $q=p^{d}$
    From the definition of $w_{0}$, $w_{1}$ in (7), if $w_{0},x_{0}>\mbox{ \rm{mod} }p=0$, $w_{1}\mbox{ \rm{mod} }p\neq 0$.
@@
    To construct a solution $x_{1}$, the component defining the solutions is chosen as $x_{1}=0$ when $w_{0},x_{0}>\mbox{ \rm{mod} }p\neq 0$ and $w_{1},x_{1}>\mbox{ \rm{mod} }\neq 0$ when $w_{0},x_{0}>\mbox{ \rm{mod} }p=0$.
Proof follows from the steps in Lemma (2) and Theorem (2) and from the fact that $x_{1}$ exists when $w_{1}\mbox{ \rm{mod} }p\neq 0$ such that $w_{1},x_{1}>\mbox{ \rm{mod} }p\neq 0$. Since there is only one prime divisor $p$ in this special case of $n=q^{r}$ and $q=p^{d}$ the construction of $x_{1}$ using CRT is not necessary as in the general case.
@@ 3.1 Decomposition modulo $p^{r}$ and solution for general modulus $n$
The Problem 1 has a solution iff for each prime divisor $p$ of $n$ the modular system of Problem 2 has a solution $x(p)$ for each $p$ and $q=p$, $r=r(p)$ such that $x(p)$ satisfies $w,x(p)>\mbox{ \rm{mod} }p\neq 0$. If $x(p)$ is a solution of the Problem 2 for each of the prime divisors $p$ of $n$ satisfying the constraint $w,x(p)>\mbox{ \rm{mod} }p\neq 0$ and
@@
If the modular system $Ax=b\mbox{ \rm{mod} }n$ has a solution $x$ then $x$ can be obtained uniquely from the residues $x(p)=x\mbox{ \rm{mod} }p^{r(p)}$ which are solutions of $Ax=b\mbox{ \rm{mod} }p^{r(p)}$ (termed as equations of residues modulo $p^{r(p)}$) and conversely if the residue equations have solutions $x(p)$ then the vector $x$ obtained by CRT with respect to coprime moduli $p^{r(p)}$ satisfies the modular system. Further the solution $x$ satisfies <w,x><w,x> coprime to $n$ iff $w,x>\mbox{ \rm{mod} }p\neq 0$ for all prime divisors $p$. Hence every solution $x(p)$ of the residue equation must satisfy $w,x>\mbox{ \rm{mod} }p=<w\mbox{ \rm{mod} }p,x(p)>\mbox{ \rm{mod} }p\neq 0$. Conversely any solutions $x(p)$ of residue systems satisfying the constraint $w,x(p)>\mbox{ \rm{mod} }p\neq 0$ determine a unique solution $x$ of Problem 2 modulo $n$. All other solutions can be obtained from such residues with respect to all prime divisors $p$. ∎
This theorem shows the decomposition of the problem from modulus $n$ to the smaller moduli $p^{r(p)}$ for prime divisors $p$ of $n$. Such a decomposition is computationally advantageous as we show next because the Smith form computation of $A,-p^{r(p)}I]$ does not involve computation of GCDs of elements of $A$ along with $n$ but only involves division of $p^{r}$ to find the largest divisor. Hence solving the decomposed linear systems modulo $p^{r(p)}$ creates a notion of relative invariant factors which are invariant factors of $A,-p^{r(p)}I]$. These can be computed much efficiently because the prime $p$ is known (hence also all the divisors of $p^{r}$) and hence analogous computation of the GCD of rows and columns of $A$ required in Smith form computation of $A$ are not required.
3.2 Smith form reduction of $A,-p^{r}I]$
The above theorem gives solution of the special case of the modular system with constraint defined in Problem 2 for $q=p$. Hence it is desirable to simplify the algorithm for computation of the invariant factors of $A,-p^{r}I]$ since all of these are powers of the prime $p$.
@@ 3.2.1 Solution of the Bezout identity with a prime power dividend
The Smith form reduction of a matrix such as $A,-p^{r}I]$ with $p$ a prime by unimodular transformations, is based on computation of integral solutions $x_{i},y$ to the Bezout identities such as
@@
d=rgd=r-g, where we now consider $a_{i}$, $x_{i}$, $y$ to be in the modular ring $\mathbb{Z}/(p^{d})$. Consider a solution in the simplest case
@@
[…]
@@
then $xb$, ybyb give solutions of the Bezout identity for right hand side $b$. The unique solutions $x_{i}$, $y_{i}$ of co-efficients of solutions $x,y$ of (10) modulo $p^{d}$ are given by the following iterations
[…]
@@
Proof follows by substitution of $p$-adic representations of $a$, $x$, $y$ in (10) and comparing co-efficients of powers of $p$. While each co-efficient is computed modulo $p$, the carry part which is quotient by $p$ is added to the higher power of $p$ shown as $t_{(k-1)}]$ in the next value of $t_{k}$ in the iteration. Note that each equation $t_{k}=0$ for $k>0$ has the unknown $x_{k}$, is of the form $a_{0}x_{k}=b_{k}\mbox{ \rm{mod} }p$ where $b_{k}$ depends on previously computed $x_{0},\ldots,x_{(k-1)}$ and only inversion modulo $p$ of $a_{0}$ is required for solving all $x_{k}$. Since $x,y$ were assumed with representations in $\mathbb{Z}/(p^{d})$ the solution $x,y$ of the iterations (12) satisfy
@@
for $x$ modulo $p^{d}$. Once $x$ is solved, $ax=1+py$ for some $y$ which can be obtained by dividing ax1ax-1 by $p^{d}$. Since $a$ is assumed to be in $\mathbb{Z}/(p^{d})$ for any other integer $a$ the Bezout identity solutions $x,y$ can be found by solving $(a\mbox{ \rm{mod} }p^{d})x=1$ modulo $p^{d}$.
@@
$a,p^{d}]Q=[1,0]$
Also $ab,p^{d}b]Q=[b,0]$. We can thus formally state
@@
If $\mbox{GCD}(a,p^{r})=b$ then there exists a unimodular matrix $Q$ such that $a,p^{r}]Q=b$.
@@
Note that the unimodular matrix $Q$ is obtained using $x,y$ the solutions of the Bezout identity (10) for the pair $(ap^{-g},p^{d})$ for d=rgd=r-g which is coprime. This computation is performed as shown in iterations (12) in Lemma 3 which involves division operations by $p$ to compute the co-efficients mod $p$ and the carry as quotient by $p$. The operations required to compute $x$, $y$ are 1) one mod $p$ inversion, 2) maximum of $d^{2}$ multiplications modulo $p$ and 3) $d$ integral divisions by $p$. The overall cost of computations of iterations (12) can be shown to be $O(l(p)d^{2})$. In general the cost of computation of the extended GCD of two integers $a,b$ is known to be approximately $O(l(a)l(b))$ where $l(a)$ is the bit length of $a$ [8]. If $a,b$ have prime factorizations
@@ 3.2.2 Solution of Bezout identity mod $p^{r}$ with multiple elements
Let $GCD(a_{1},a_{2},\ldots,a_{N},p^{r})=p^{g}$. Then there exists a unimodular matrix $Q$ such that $a_{1},a_{2},\ldots,a_{n},p^{r}]Q=[p^{g},0,\ldots,0]$. The first column of $Q$ is $x_{1},x_{2},\ldots,x_{N},y]^{T}$ where $x_{i}$ and $y$ are modulo $p^{r}$ and satsfy
@@
Let $Q_{1}$ be a $2$ unimodular matrix as computed in Lemma 3 such that $a_{N},p^{r}]=[p^{t_{1}},0]$ where $p^{t_{1}}$ is the GCD of $(a_{N},p^{r})$. By induction step assume that the GCD of $(a_{(N-1)},\ldots,a_{1},p^{r})$ be $p^{t_{(N-1)}}$ and let $Q_{(N-1)}$ be a $n\times n$ unimodular matrix such that
$a_{(N-1)},\ldots,a_{1},p^{r}]Q=[p^{(t_{(N-1)}},0,\ldots,0]$
The GCD of $(a_{N},p^{t_{(N-1)}})$ is then $p^{g}$. Let $Q^{\prime}$ be a unimodular matrix as derived in lemma 4 such that $a_{N},p^{t_{(N-1)}}Q^{\prime}=[p^{g},0]$. Define
@@
The construction of unimodular matrix for solving the Bezout identity modulo $p^{r}$ for multiple elements $(a_{1},\ldots,a_{N})$ and GCD $p^{g}$ with $p^{r}$ uses the computations of Lemma 3 which are mod $p$ and quotients by $p$. Hence for a known prime $p$ these computations are much reduced compared to GCD computation of general $(N+1)$ elements and further solution of the general Bezout identity. In the next proposition we derive the unmodular matrix to transform the matrix $A,-p^{r}I]$ to its Smith form.
@@
First bring the matrix $\hat{A}=[A,-p^{r}I]$ by column (unimodular) transformations $\hat{A}V_{1}$ to the upper triangular form $p^{r}I,A]$. Let $t\leq r$ be the largest power such that $p^{t}$ divides all elements in the first row of $A$. Then $p^{t}$ is the GCD of elements in the first row of $\hat{A}V_{1}$. Hence there exists a unimodular matrix $V_{2}$ obtained from the extended Euclidean divison representation of the first row such that $\hat{A}_{2}=\hat{A}V_{1}V_{2}$ has $p^{t}$ as $(1,1)$ element while the entire first row is zero. Repeat the same process to make the first column of $\hat{A}_{2}$ equal to zero by constructing a unimodular matrix $P$ and computing $P\hat{A}_{2}$. Repeat both steps of operation by left and right unimodular matrices till $\hat{A}$ is brought to the form
@@
Since the original matrix $\hat{A}$ has some of its $2\times 2$ minors of the form $p^{2r}$ the GCD of elements of $A_{1}$ must have a divisor $p^{k_{2}}$ where $k_{2}\geq k_{1}$. Now repeat the same steps to make $A_{1}$ successively into diagonal form by unimodular row and column operations. Hence after repeating such steps the matrix $A,-p^{r}]$ is transformed into a diagonal form with diagonal entries $p^{k_{i}}$. Next add all columns from 2nd to last to the first column and then reduce the matrix to $p^{r_{1}}$ as $(1,1)$ entry and and another matrix $A_{1}$ as above. Now we can see that $0\leq r_{1}$ and $p^{r_{1}}$ divides all other powers $p^{k_{i}}$ for $i\geq 2$. Repeating the same process now on the resultant matrix gives the Smith form of $A,-p^{r}I]$ as claimed. ∎
@@
The reduction to Smith form of $A,-p^{r}I]$ shown in the above Proposition uses division by $p$ in all the steps to determine largest powers of $p$ as the gcd of entries of the matrices at each step. If any element in the matrix is not divisible by $p$ then the corresponding invariant factor is $1$. Using the well known result on high probability of coprimeness of random integers it follows that the invariant factors for lower indices are highly likely to be $1$.
@@
Computing the GCD and solving the Bezout identity of a set of $N$ integers and $p^{r}$ modulo $p^{r}$ has complexity equivalent to the $O(Nd^{2}(\log p)^{2})$ where d=rtd=r-t and $p^{t}$ is the GCD of $(a_{1},\ldots,a_{N},p^{r})$. For a given prime $p$ the unimodular matrix $Q$ which can transform $a_{1},a_{2},\ldots,a_{N},p^{r}]$ as
@@ 3.3 High level algorithm to solve the modular system
4:  Form the augmented matrix $A,-p^{r}I]$.
5:  Find unimodular integral matrices $P$, $Q$ from the procedure described in Lemma 2, to determine the Smith form $S_{p},0]$ of the augmented matrix.